Red Acre Industries paid a dividend of $0.63 per share after having earnings of $2.45, and has an ROE of 8 percent. What is Red Acre Industries' payout ratio? retention ratio? expected growth rate? please show steps, it would be helpful to me. thanks!
Answer #1

payout ratio=Dividend/earnings

=(0.63/2.45)=0.2571(or 25.71% approx).

Retention ratio=1-payout ratio

=(1-0.2571)=0.7429(or 74.29% approx).

Growth rate=ROE*Retention ratio

=(0.08*0.7429)=5.94%(Approx).